Find the volume of the solid inside the sphere x^2+y^2+z^2=9 and inside the x^2+y^2=1clinder
1
Expert's answer
2020-12-25T12:36:55-0500

"x^2+y^2+z^2=9"

and

"x^2+y^2=1"

Using

"r^2 =x^2 +y^2"

"r^2 =1"

"z=\sqrt{ (9-r^2)}"

The volume of the solid in cylindrical coordinates is:


"V= \int_{0}^{2\pi} \int_{0}^{1} \int_{-\sqrt{9-r^2}}^{\sqrt{9-r^2}} rdzdrdθ"


Evaluating the first integral, we have:

"V= \int_{0}^{2\pi} \int_{0}^{1}2r \sqrt{9-r^2}drdθ" ...


Evaluating the second integral, we have:


"V= \int_{0}^{2\pi} 18 - \frac {32\sqrt2}{3}dθ"


Evaluating in terms of the angle, we have;

"V= 18(2\pi) - \frac {32\sqrt2}{3} (2\pi)"


"V= 36\pi - \frac {64 \pi\sqrt2}{3}"